# NOLO: Navigate Only Look Once

### v1

@@ Abstract
The in-context learning ability of Transformer models has brought new possibilities to visual navigation. In this paper, we focus on the video navigation setting, where an in-context navigation policy needs to be learned purely from videos in an offline manner, without access to the actual environment. For this setting, we propose Navigate Only Look Once (NOLO), a method for learning a navigation policy that possesses the in-context ability and adapts to new scenes by taking corresponding context videos as input without finetuning or re-training. To enable learning from videos, we first propose a pseudo action labeling procedure using optical flow to recover the action label from egocentric videos. Then, offline reinforcement learning is applied to learn the navigation policy. Through extensive experiments on different scenes, we show that our algorithm outperforms baselines by a large margin, which demonstrates the in-context learning ability of the learned policy.
@@ 1 Introduction
Visual navigation has long become a research focus due to its wide application in different fields, including mobile robots [2], autonomous vehicles [3], and virtual assistants [4]. A large body of research has been proposed to solve this problem via either combining modules like object detectors, key point matchers, depth estimators, and SLAM modules [5, 6, 7, 8, 9] or training a single policy model end-to-end [1, 10, 11]. Although these methods have achieved great success, they often suffer from poor generalization or require additional interactions and training when transferred to a novel scene, which leads to non-trivial problems in real-world applications.
Unlike these methods, humans can solve the visual navigation problem, more specifically the indoor visual navigation problem, in a much easier way. Instead of walking around and exploring a new room in person, we possess an innate ability to easily understand the room configuration by just watching an egocentric traversal video of this room and then figure out how to reach different targets in the room. However, unlike humans, the majority of existing methods heavily rely on heterogeneous sensors to perceive the environment, such as pose sensors like GPS+Compass [12, 13] for estimating position and orientation, RGB-D cameras [14] for observing the surrounding environment and measuring the distance, and IMU [15] for tracing movement. Whereas humans, do not explicitly rely on poses or distance measurement and can still navigate to different targets.
Inspired by human navigation, a natural and intuitive idea would be offline training a navigation policy that can generalize to a new scene given an egocentric traversal video of the new room, which aligns with the idea of in-context learning [16]. It has been shown that large language models (LLMs) can work well on a new task when few demonstrations of the corresponding task are given as the context, without any additional training process [17]. The in-context learning method has proven its value in many fields ranging from LLM-based agents [18], to meta-algorithm learning [19, 20, 21]. Given the success of in-context learning, we propose a new setting of the visual navigation problem, Video Navigation. The goal of video navigation is to train an in-context policy to find objects occurred in the context video. During the training process, we have only access to a dataset of egocentric traversal videos in different scenes and we need to train an in-context navigation policy in an offline manner. During inference, the video of a new scene will be given as the context, and the learned policy should take the context video as input and generalize to the new scene without finetuning or re-training.
The setting of video navigation is promising to equip mobile agents with human-like navigation capabilities. With video navigation, imagine that a new sweeping robot can quickly and freely navigate to any corner of your house and do cleaning after only watching a short video clip of the house layout. Thus, we can say video navigation is more suitable for real-world applications, as finetuning or re-training the policy is usually infeasible after deployment. Although encouraging, video navigation poses considerable challenges:
@@
  • First, due to the lack of sufficient information, including actual actions, spatial correlation, and visual understanding of complex scenes, learning to navigate by only watching egocentric traversal videos is challenging.
  • Second, egocentric traversal videos do not exhibit explicit intentions. As a result, inverse reinforcement learning or learning from demonstrations cannot be directly applied, since expert trajectory is an imperative prerequisite.
@@
  • More importantly, in the video navigation setting, we have no access to camera intrinsics, robot poses, maps, odometers, and depth inputs, which distinguishes our work from others. Most existing methods cannot leave with poses or depth images to explicitly build a map for further planning. In this work, we expect the agent can navigate like a human, only with egocentric observations.
To address these challenges, we propose Navigate Only Look Once (NOLO). NOLO learns a generalizable in-context navigation policy conditioned on a context video, such that it can adapt to the new scene without finetuning or re-training by understanding the corresponding context video. To train such a policy, we first extract pseudo actions from context videos using optical flow to get complete context trajectories. Subsequently, an offline reinforcement learning method is adopted to learn the in-context navigation policy. Our key contributions can be summarized as follows:
@@
  • We present Video Navigation, a novel and practical setting for visual navigation. To the best of our knowledge, it is the first attempt to learn an in-context navigation policy in an offline manner purely from videos such that the learned policy can adapt to different scenes by taking corresponding videos as context.
@@
  • Empirical evaluations in RoboTHOR and Habitat demonstrate the effectiveness of NOLO in video navigation. Notably, this is achieved by only watching a single 30-second video clip in each scene.
@@ 2 Related Work
Visual Navigation. The visual navigation problem can be categorized into indoor navigation [22, 23, 24] and outdoor navigation [25, 26]. In this paper, we only focus on the indoor navigation problem, such that all scenes are indoor scenarios. For indoor navigation, research mainly focuses on finding the location of specific positions [13, 27], rooms [28], or object instances [29, 9, 27, 30]. The goals of navigation can be described in coordinates [13], natural language [31, 8, 32, 11, 33], goal image of target position [28, 8, 32], or goal image of object instances [29, 9, 27, 30]. Our video navigation of finding objects occurred in the context video falls into the last group.
Visual navigation methods can also be categorized by whether exploration is needed [34]. Most works pursue generalization over unseen goals or unseen environments and usually rely on learning algorithms that require massive online interactions, e.g., reinforcement learning [35, 36, 37], to explore the environment and accomplish the task. While some works try to improve data efficiency, these methods still need to interact with the environment using a random policy to learn an inverse dynamic model [38] or assume high-quality datasets including features like robot poses [7] or depth images [14]. Unlike previous work, our method frees robots from interacting with actual environments and targets to only learn from videos without actual actions.
In-Context Learning. The ability of in-context learning is first demonstrated in large language models [17, 39, 40], where the learned model can perform a new task, given a few corresponding examples as context. Similar abilities are also observed in vision language models [41, 42, 43] and models for other modalities [44, 45]. The ability of in-context learning can be applied to AI agents [46, 47] to solve different tasks. It has also been shown that large Transformer models can be used to learn a meta-algorithm to imitate the reinforcement learning algorithm [19, 20, 48]. Our work focuses on the fusion of in-context learning and the visual navigation problem, and to the best of our knowledge, is the first attempt to train an in-context navigation policy that can adapt to different scenes by taking corresponding videos as context.
@@
Learning from Videos. There are four main lines of research study learning from videos. (1) Learn distinctive representations for intrinsic rewards. These approaches mainly include predicting the future [49, 50, 51], learning temporal abstractions [51, 52, 53, 54], contrastive learning [55, 56], adversarial learning [57, 58, 59, 60] or other transition-learning based methods [61, 62, 63]. (2) Recover value function along with representations [64, 65, 66, 67]. (3) Pretrain action-free model in the first stage and then boost the model with additional action information [68, 69, 70]. (4) Discover latent actions [71, 72, 73, 74] and then align to the real actions online in the environment [71, 72, 73], by human annotation [74] or from offline datasets [73]. Few attempts are devoted to decoding real actions directly from videos. A recent work [75] attempts to principally use similar rule-based methods to directly infer actions for robot manipulation. However, it does not exhibit in-context learning capabilities and requires additional depth image inputs. Unlike previous work, we focus on the visual navigation problem and by utilizing its specificity, propose an offline method that directly learns the actual navigation policy, not the latent one, from videos.
@@ 3.1 The Problem Formulation of Video Navigation
In video navigation, we offline recover a navigation policy $\pi_{\theta}\left(a|g,o,\mathcal{V}\right)$, which takes a goal image $g$, current observation $o$, and context video $\mathcal{V}$ as input and produces action $a$ as output. Such a policy is expected to find objects occurred in the context video. To train such a policy, we assume a video dataset $\mathcal{D}$ contains $N$ videos of $N$ different scenes, $\mathcal{D}=\left\{\mathcal{V}^{1},\mathcal{V}^{2},...,\mathcal{V}^{N}\right\}$. Each video $\mathcal{V}^{i}$ contains $T$ frames $f^{i}_{t}$, $\mathcal{V}^{i}=\left\{f^{i}_{1},f^{i}_{2},...,f^{i}_{T}\right\}$. During inference, we deploy the learned policy in a new scene given a context video $\mathcal{V^{\textrm{new}}}\notin\mathcal{D}$ and a goal frame $g\in\mathcal{V^{\textrm{new}}}$. The policy is expected to adapt to this new scene via the context video without any finetuning or re-training.
@@ 3.2 Pseudo Action Labeling
To recover a policy from videos, lots of work adheres to the paradigm of latent action extraction from adjacent frames utilizing an inverse dynamics model (IDM) [76, 71, 72, 73, 74], followed by latent policy learning and policy decoding using actual action labels. Within the context of visual navigation, we propose to use GMFlow [77] as an implicit IDM, $F_{\xi}$, to estimate discrete pixel displacements between two consecutive frames, thereby enabling the prediction of the navigation action $\hat{a}_{t}$ with simple rule-based post-processing $p$. Continuous actions like linear velocity and angle velocity are usually applied to real-world navigation, while high-level discrete actions are also common in competitions like habitat navigation challenge [78] and modern navigation tasks [25]. The sequential process can be formalized in Equation (1),
$\Delta f_{t}$ $=F_{\xi}(f_{t},f_{t+1})$ (1)
$\hat{a}_{t}$ $=p(\Delta f_{t}).$
Occasionally, we notice that the displacement of pixels observed across frames manifests irregular patterns. To mitigate this issue, for each pair of adjacent frames $(f_{t},f_{t+1})$, we filter dominant vectors $\nu_{t}$ that exhibit gradient amplitudes within the upper decile, which are subsequently discretized into actions derived from the action space, making it possible to trace the trajectory of the agent over time. The pseudo action labeling is illustrated in Figure 2, and see Appendix B.1 for more details.
@@ 3.3 In-Context Policy Modeling
$h_{t+1},\pi_{\theta}\left({a}_{t}|g,o_{t},\mathcal{T}\right)=\text{VN}\circlearrowleft\text{Bert}(g,o_{t},\mathcal{T}).$ (2)
@@
At the initialization stage, a zero-padded hidden state $h_{\text{init}}$ is additionally inserted at the back of the context embedding sequence for recurrence. The entire token embedding sequence $\boldsymbol{e}_{\text{init}}=[h_{\text{init}},e^{s}_{\text{init},1},e^{a}_{\text{init},1}\dots,e^{a}_{\text{init},T-1},e^{s}_{\text{init},T}]$ is encoded by a multi-layer bidirectional self-attention module (SA) as $\boldsymbol{e}_{0}=[\boldsymbol{e}^{c},h_{0}]=\text{SA}(\boldsymbol{e}_{\text{init}})$. The output feature of the hidden state $h_{0}$ serves as the initial compressed representation of the context trajectory. After initialization, the context embedding sequence $\boldsymbol{e}^{c}$ keeps unchanged. At each timestep $t$, the current observation $o_{t}$ and the goal image $g$ are encoded into $e^{s}_{t}$ and $e^{g}$, respectively. The concatenated embedding sequence $\boldsymbol{e}_{t}=[h_{t},\boldsymbol{e}^{c},e^{s}_{t},e^{g}]$ is then fed into the cross-attention module (CA) to obtain the updated embedding $\boldsymbol{e}_{t+1}=\text{CA}(\boldsymbol{e}_{t})$ which is further aggregated and projected to action distribution $\pi_{\theta}\left({a}_{t}|g,o_{t},\mathcal{T}\right)$ and the next hidden state $h_{t+1}$. The updated hidden state $h_{t+1}$ is subsequently processed by a multilayer perceptron (MLP) $P_{Q}$ to regress Q-values $Q_{\omega}\left(g,o_{t},a_{t},\mathcal{T}\right)$ (discussed later in Section 3.4) and by another MLP $P_{D}$ to predict the binary terminal signal $\delta_{\upsilon}\left(d_{t}|g,o_{t},\mathcal{T}\right)$, where $d_{t}$ is used for predicting the termination of the current task.
@@ 3.4 Batch-Constrained Q-Learning
As the policy used to record the video could be highly suboptimal for the video navigation problem, we prefer offline reinforcement learning over imitation learning for policy training. Specifically, we adopt BCQ [81], an offline reinforcement learning method that emphasizes constraining the action selection to those within the distribution of the observed frame-action pairs. The Bellman update in Equation (4) is used for learning Q-function $Q_{\omega}$, given the temporal difference target in Equation (5) where $r(g,o_{t},\hat{a}_{t})=\mathbbm{I}(o_{t}=g)$ is a binary reward indicating the success of a task,
$\mathcal{L}_{q}=\mathop{\mathbb{E}}_{o_{t},o_{t+1},g,\hat{a}_{t}\sim\mathcal{T}^{i}}\left[Q_{\omega}\left(g,o_{t},\hat{a}_{t},\mathcal{T}^{i}\right)-y_{t}\right]^{2},$ (4)
$y_{t}=r(g,o_{t},\hat{a}_{t})+\gamma\max_{\tilde{a}_{t+1}\in A^{\beta}}Q_{\omega}\left(g,o_{t+1},\tilde{a}_{t+1},\mathcal{T}^{i}\right).$ (5)
Since the action space is discrete, to select $\tilde{a}_{t+1}$, we define a set $A^{\beta}$, where $\beta$ denotes the threshold for the action selection. We restrict the action choice considering the action distribution in the context trajectory dataset,
@@
The learned Q-function $Q_{\omega}$ can guide the agent towards the given goal $g$.
@@ 3.5 Temporal Coherence
Learning temporally aligned representation of videos is crucial for understanding the causality of the context video. Inspired by RLHF [82], we learn a "fuzzy" measure of progress from their natural temporal ordering, where the higher preference score is assigned to the later frames in the video. To achieve this, we learn a temporal indicator $\hat{u}_{\phi}(e)$, which consists of 1D self-attention and spectral normalization modules to map a visual context embedding $e^{s}_{t}$ to a utility score. Such an indicator can be optimized as follows:
$\min_{\zeta,\phi}\mathcal{L}_{t}=-\mathbb{E}_{f_{t_{-}},f_{t_{+}}\sim\mathcal{V}}\log\sigma\left[\hat{u}_{\phi}\left(E_{\zeta}(f_{t_{+}})\right)-\hat{u}_{\phi}\left(E_{\zeta}(f_{t_{-}})\right)\right],$ (7)
where $t_{+}>t_{-}$ and $f_{t_{+}},f_{t_{-}}$ stands for the later and earlier frame in the video. This self-supervised paradigm naturally drives the temporal indicator $\hat{u}_{\phi}(s)$ to maximize likelihood to ensure the temporal order preference $f_{t_{+}}\succ f_{t_{-}}$, automatically leading to temporally aligned context visual representation, which is beneficial to better understand the temporal relation in a novel scene, as demonstrated later in Section 4.5.

### v2

@@ Abstract
The in-context learning ability of Transformer models has brought new possibilities to visual navigation. In this paper, we focus on the video navigation setting, where an in-context navigation policy needs to be learned purely from videos in an offline manner, without access to the actual environment. For this setting, we propose Navigate Only Look Once (NOLO), a method for learning a navigation policy that possesses the in-context ability and adapts to new scenes by taking corresponding context videos as input without finetuning or re-training. To enable learning from videos, we first propose a pseudo action labeling procedure using optical flow to recover the action label from egocentric videos. Then, offline reinforcement learning is applied to learn the navigation policy. Through extensive experiments on different scenes both in simulation and the real world, we show that our algorithm outperforms baselines by a large margin, which demonstrates the in-context learning ability of the learned policy. For videos and more information, visit our project page.
@@ 1 Introduction
Visual navigation has long become a research focus due to its wide application in different fields, including mobile robots [18], autonomous vehicles [38], and virtual assistants [22]. Existing approaches to visual navigation can be broadly divided into two main categories: modular and end-to-end methods. Modular methods decompose the navigation problem into specific tasks, such as simultaneous localization and mapping, object segmentation, keypoint matching, and depth estimation [15, 16, 14, 91, 45]. End-to-end methods [37, 77, 10], on the other hand, bypass such modularization, directly learning the mappings from raw sensory observations (e.g., GPS and Compass [31, 65], RGB-D cameras [34], and IMU [58]) to action outputs. While these approaches have yielded promising results, they still face significant challenges, particularly in terms of generalizability. Modular methods can be brittle when transferred to unfamiliar environments. End-to-end approaches mainly focus on online learning [16, 88, 17], which requires continuous, interactive engagement with simulated or real environments. Thus, they may require additional interactions or exploration to adapt to novel scenes, complicating their deployment in real-world navigation settings.
To address these limitations, it is helpful to consider how humans intuitively navigate. We can often find our way through unfamiliar spaces after watching a simple traversal video. Our navigation abilities do not depend on precise positional data or distance measurement. Instead, we rely primarily on visual recognition, understanding, and memory. Building on this insight, we introduce a new visual navigation setting called Video Navigation, which mirrors how humans rely on visual observations to navigate. In video navigation, a mobile agent is provided with a context video of an unfamiliar environment and learns to reach any target seen in the video. For instance, a sweeping robot equipped with video navigation capabilities could immediately navigate and clean any area within a house after viewing a short video of its layout. We believe video navigation has immense potential for endowing mobile agents with human-like navigation capabilities. Although encouraging, achieving robust video navigation poses considerable challenges:
@@
  • Limited Observations. Learning to navigate from egocentric traversal videos is challenging due to a lack of information, including actual actions, spatial cues, scene structure, and visual context.
  • Absence of Explicit Intent. Egocentric videos lack clear navigational intentions, making conventional methods like inverse reinforcement learning unsuitable, as these approaches typically require expert trajectory data.
@@
  • Minimal Sensory Input. Camera intrinsics, robot poses, maps, odometers, or depth inputs are not necessary for video navigation, which distinguishes our work from others. A majority of existing methods prefer to build explicit spatial maps for planning with heterogeneous sensors. In this work, we aim to enable human-like navigation using only egocentric RGB observations.
Fundamentally, the objective of video navigation is to train a meta-policy that enables a robot to adapt to a novel scene–essentially, to navigate anywhere using cues from a context video. Recent research demonstrates that in-context learning (ICL [23]) excels at solving tasks using a few contextual demonstrations without requiring additional training [7] and has proven valuable across fields from LLM-based agents [85] to meta-learning [48, 49, 55]. Inspired by the success of ICL, we propose Navigate Only Look Once (NOLO), which learns a transferable in-context navigation policy conditioned on a context video, enabling adaptation to new scenes without fine-tuning or further re-training.
To achieve this, we collect a dataset of egocentric traversal videos, each covering a unique scene. We derive pseudo-actions from these context videos using optical flow, generating full frame-action trajectories, and then employ an offline reinforcement learning approach to train the navigation policy. This resource-efficient offline training pipeline eliminates the need for time-consuming online interactions, enabling large-scale training on simulated data and facilitating the transfer of navigation skills to unfamiliar environments. Furthermore, NOLO’s in-context adaptability makes it particularly suitable for real-world deployment, where frequent policy adjustments are often impractical.
Our key contributions can be summarized as follows:
@@
  • We present a novel and practical setting Video Navigation, which necessitates learning an in-context navigation policy in an offline manner purely from videos such that the learned policy can adapt to different scenes by taking corresponding videos as context.
@@
  • Empirical evaluations on RoboTHOR and Habitat benchmarks demonstrate the advantages of NOLO over baselines in video navigation. Notably, navigation is achieved by only viewing a single 30-second video clip per scene. We further deploy NOLO on the Unitree Go2 robot and experiments verify that NOLO is also effective and friendly to real-world deployment.
@@ 2 Related Work
Visual Navigation. The visual navigation problem can be categorized into indoor navigation [11, 33, 93, 30, 54, 61, 46, 77, 60, 42, 51] and outdoor navigation [52, 76, 86]. In this paper, we only focus on the indoor navigation problem. Current research mainly focuses on finding the location of specific positions [65, 41], rooms [62], or object instances [44, 45, 41, 50]. The goals of navigation can be described in coordinates [65], natural language [68, 91, 90, 10, 94], goal image of target position [62, 91, 90], or goal image of object instances [44, 45, 41, 50, 43, 60, 46]. Our video navigation of finding objects occurred in the context video falls into the last group. Visual navigation methods can also be categorized by whether exploration is required. Most existing works aim for generalization across new goals or environments, heavily relying on learning algorithms that require extensive online interactions, often through reinforcement learning [99, 80, 2], paired with real-time visual feedback [61]. This approach enables the agent to adapt dynamically but requires significant interaction within the environment. Some studies attempt to improve data efficiency but still rely on random policy exploration or assume access to rich datasets, including robot poses [14] or depth images [34, 30, 42, 45]. In contrast, NOLO learns entirely offline. It leverages video data for navigation, freeing the agent from real-world exploration and allowing effective policy learning without performing actual actions, marking a novel approach in video-based visual navigation.
@@
Learning from Videos. There are four main lines of research study learning from videos. (1) Learn distinctive representations for intrinsic rewards. These approaches mainly include predicting the future [98, 25, 96], learning temporal abstractions [96, 75, 4, 8], contrastive learning [47, 79], adversarial learning [66, 82, 40, 57] or other transition-learning based methods [6, 53, 64]. (2) Recover value function along with representations [26, 59, 32, 5]. (3) Pretrain action-free model in the first stage and then boost the model with additional action information [92, 74, 87]. (4) Discover latent actions [24, 56, 72, 9] and then align to the real actions online in the environment [24, 56, 72], by human annotation [9] or from offline datasets [72]. Unlike previous work, we focus on the visual navigation problem and by utilizing its specificity, propose an offline method that directly learns the actual navigation policy, not the latent one, from videos. Few attempts are devoted to decoding real actions directly from videos. A recent work AVDC [43] attempts to principally use similar rule-based methods to directly infer actions for robot manipulation. However, it is an online algorithm with additional depth image inputs. We also compare NOLO with AVDC [43] in Sec. 4.5.
In-Context Learning. In-context learning (ICL), first demonstrated in large language models [7, 83, 19], enables models to perform new tasks with a few examples as context. This capability has extended to vision-language models [3, 84, 1], AI agents [78, 36], and models across heterogeneous modalities [39, 27]. Our work uniquely applies ICL to video navigation. To our knowledge, NOLO is the first attempt to learn an adaptable policy guided by context videos for diverse scenes.
@@ 3.1 The Problem Formulation of Video Navigation
In video navigation, we offline recover a navigation policy $\pi_{\theta}\left(a|g,o,\mathcal{V}\right)$ to find objects occurred in the context video, which takes a goal image $g$, current observation $o$, and context video $\mathcal{V}$ as input and outputs discrete action $a$, considering high-level discrete actions are commonly used in competitions like the Habitat Navigation Challenge [69] and modern navigation tasks [52]. To train such a policy, we assume a video dataset $\mathcal{D}$ contains $N$ videos collected (e.g., by some simple rule-based policy or humans) in $N$ different scenes, $\mathcal{D}=\left\{\mathcal{V}^{1},\mathcal{V}^{2},...,\mathcal{V}^{N}\right\}$. Each video $\mathcal{V}^{i}$ contains $T$ frames $f^{i}_{t}$, $\mathcal{V}^{i}=\left\{f^{i}_{1},f^{i}_{2},...,f^{i}_{T}\right\}$. During inference, we deploy the learned policy in a new scene given a context video $\mathcal{V^{\textrm{new}}}\notin\mathcal{D}$ and a goal frame $g\in\mathcal{G}^{\textrm{new}}$ describing an object occurred in $\mathcal{V^{\textrm{new}}}$. The policy is expected to adapt to this new scene via in-context learning.
@@ 3.2 Pseudo Action Labeling
In video navigation, capturing environmental dynamics is crucial for effective adaptation to new scenes. Building on prior work [29, 24, 56, 72, 9] that leverages inverse dynamics models (IDMs) to extract latent actions from adjacent frames, we make a further step by converting latent actions into actual discrete actions within video navigation. Specifically, we propose pseudo action labeling, as illustrated in  Fig. 2. we employ GMFlow[89] as an implicit IDM $F_{\xi}$ to estimate discrete pixel displacements between consecutive frames and predict real navigation actions from flow maps with rule-based post-processing. For each context video $\mathcal{V}$, we label a sequence of pseudo-actions $\boldsymbol{\hat{a}}$ to get a trajectory $\mathcal{T}=\{f_{t},\hat{a}_{t}\}^{T}_{t=1}$. See Appendix 6.1 for more details.
@@ 3.3 In-Context Policy Modeling
$h_{t+1},\pi_{\theta}\left({a}_{t}|g,o_{t},\mathcal{T}\right)=\text{VN}\circlearrowleft\text{Bert}(g,o_{t},\mathcal{T}),$ (1)
where $h$ is specially designed as a recurrent hidden state to capture informative attention history.
@@ 3.4 Batch-Constrained Q-Learning
Since the policy used to record the video may be highly suboptimal for video navigation, we favor offline reinforcement learning over imitation learning for policy learning. Specifically, we adopt BCQ [28], an offline reinforcement learning method that emphasizes constraining the action selection to those within the distribution of the observed frame-action pairs. Q-function $Q_{\omega}$ is learned from Bellman update in Eq. 3, where $r(g,o_{t},\hat{a}_{t})=\mathbbm{I}(o_{t}=g)$ is a binary reward indicating the success of a task,
$\begin{split}\mathcal{L}_{q}=\mathop{\mathbb{E}}_{o_{t},o_{t+1},g,\hat{a}_{t}\sim\mathcal{T}^{i}}\left[Q_{\omega}\left(g,o_{t},\hat{a}_{t},\mathcal{T}^{i}\right)-y_{t}\right]^{2},\\ y_{t}=r(g,o_{t},\hat{a}_{t})+\gamma\max_{\tilde{a}_{t+1}\in A^{\beta}}Q_{\omega}\left(g,o_{t+1},\tilde{a}_{t+1},\mathcal{T}^{i}\right).\end{split}$ (3)
Given the action space is discrete, we define a set $A^{\beta}$ in Eq. 4 to select $\tilde{a}_{t+1}$, where $\beta$ represents the threshold for action selection. Action choices are constrained based on the action distribution within the context trajectory dataset. The learned Q-function, $Q_{\omega}$, thus can wisely guides the agent toward the target goal $g$.
@@ 3.5 Temporal Coherence
Learning temporally aligned representation of videos is crucial for understanding the causality of the context video. Inspired by RLHF [63], we learn a ”fuzzy” measure of progress from their natural temporal ordering, where the higher preference score is assigned to the later frames in the video. To achieve this, we learn a temporal indicator $\hat{u}_{\phi}(e)$, which consists of 1D self-attention and spectral normalization modules to map a visual context embedding $e^{s}_{t}$ to a utility score. $\hat{u}_{\phi}$ is optimized according to Eq. 5:
$\min_{\zeta,\phi}\mathcal{L}_{t}=\underset{f_{t_{-}},f_{t_{+}}\sim\mathcal{V}}{\mathbb{E}}-\log\sigma\left[\hat{u}_{\phi}\left(E_{\zeta}(f_{t_{+}})\right)-\hat{u}_{\phi}\left(E_{\zeta}(f_{t_{-}})\right)\right],$ (5)
where $t_{+}>t_{-}$ and $f_{t_{+}},f_{t_{-}}$ stands for the later and earlier frame in the video. This self-supervised paradigm naturally drives the temporal indicator $\hat{u}_{\phi}(s)$ to maximize likelihood to ensure the temporal order preference $f_{t_{+}}\succ f_{t_{-}}$, leading to temporally aligned context visual representation, which is beneficial to better understand the temporal relation in a novel scene, as elaborated in later Sec. 4.5.